OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $78 for 1 adult ticket and 2 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 78 = one-halfx + (x - 6) What is the regular price of a child's ticket? help?

OpenStudy (anonymous):

the choices are: A. 50 b. 56 c. 75 d. 81

OpenStudy (johnweldon1993):

so we have \[\frac{ 1 }{ 2 }x + x - 6 = 78\] We combine like terms....what is 1/2x + x? \[\frac{ 3 }{ 2 }x - 6 = 78\] Add 6 to both sides \[\frac{ 3 }{ 2 }x = 84\] And finally multiply both sides by 2/3 \[x = 84 \times \frac{ 2 }{ 3 }\] so 84 times 2 = 168. and 168 divided by 3 = 56 dollars for a child's admission ticket Does that make sense?

OpenStudy (anonymous):

ohhhhhhh got the same answer! thanks! mind if i ask you one more?

OpenStudy (anonymous):

Solve the following for a and c, the regular price of an adult and child respectively.\[\left\{\frac{a}{2}+2 \frac{c}{2}=78,a-c=6\right\}\]\[\{a=56,c=50\} \]

OpenStudy (anonymous):

Solve the equation 2 over 3 k + 6 = −12 for k. Please show your work and solution in the space provided. mind helping with one more?

OpenStudy (johnweldon1993):

\[\frac{ 2 }{ 3k + 6 } = -12\] like that?

OpenStudy (anonymous):

i got that toobut i feel like its to easy to be right haha

OpenStudy (johnweldon1993):

No that is just me asking if that is the correct way the question is written?

OpenStudy (anonymous):

and also how do u do the divided sign like that on a computer?

OpenStudy (anonymous):

yea its the right set up

OpenStudy (johnweldon1993):

Okay...so multiply both sides of the equation by 3k + 6 \[2 = -12(3k + 6)\] Now distribute that -12 into the parenthesis...what do you get after that?

OpenStudy (anonymous):

2 = -36k - 72

OpenStudy (johnweldon1993):

Right....so now lets add 72 to both sides to get \[-36k = 74\] So what does 'k' equal?

OpenStudy (anonymous):

k = 2.05

OpenStudy (johnweldon1993):

Well I would keep it in fraction form \[k = -\frac{ 37 }{ 18 }\] is what I get

OpenStudy (anonymous):

why would i keep it in fraction form?

OpenStudy (johnweldon1993):

Just makes it easier to write... and think...when you turn this into a decimal....do you write 2.05? or 2.055? or 2.06 for rounding.....it's more exact in fraction form